OpenStudy (anonymous):

Find g(1) if g(x) = x^2 + 1. 2 3 4

OpenStudy (anonymous):

Insert 1 into your equation/function\[g(x) = x^2 + 1\]\[g(1) = 1^2 + 1\]What do you think is the answer?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

YUP!